OpenStudy (anonymous):

Cos 3A = 4 cos^3 A - 3 cosA

OpenStudy (anonymous):

cos (3A) = cos(2A + A) = cos (2A) cos (A) - sin(2A) sin(A) = [ 2cos^2(A) - 1 ] cos (A) - (2 sin A cos A )sin A = 2cos^3(A) - cos A - 2sin^2(A) cos A = 2cos^3(A) - cos A - 2( 1 - cos^2(A)) cos A = 2cos^3(A) - cos A - 2cos A + 2cos^3(A) = 4cos^3(A) - 3cos A

OpenStudy (anonymous):

wao, that is fast! thanks